120mL of water is added to 30mL of the 40% ethanol solution. What is the dilution factor? And, what is the concentration of the diluted solution?


1
Expert's answer
2021-03-17T08:46:42-0400

1) Dilution factor refers to the ratio of the volume of the initial (concentrated) solution to the volume of the final (dilute) solution1, that is, the ratio of V1 to V2 or V1 : V2. A dilution factor, DF, can be calculated: DF = V2 / V1.

DF = 150/30 = 5

2) Dilution factor may also be expressed as the ratio of the concentration of stock solution (C1) to the concentration of the final diluted solution (C2). Stock Solution Concentration (C1) is the concentration of the stock solution.

DF = C1 / C2

C2 = C1 / DF = 40 / 5 = 8%